Exercise 1 (2 pts). Alice has her RSA public key (n,e) where n = 247 = 13·19 and e = 25.
What is her secret key and what is her signature corresponding to the message M = 63?


n=247=13 x 19
Answer #1

n = 247 = 13 * 19
so,
p = 13 & q = 19

\phi ( n ) = (13-1) * (19 -1) = 216

e =25

secret key is d such that

e*d = 1 mod 216

using eucledian theorem

d is multiplicative inverse of e in Z216

q r1 r2 r t1 t2 t
8 216 25 16 0 1 -8
1 25 16 9 1 -8 9
1 16 9 7 -8 9 -17
1 9 7 2 9 -17 26
3 7 2 1 -17 26 -95
2 2 1 0 26 -95 216
1 0 -95 216

t = t1 -q*t2

t1 = t2

t2 = t

value of t = -95 or (216 - 95) = 121

so d =121

where secret key = (d,n) = (121 , 247)

signature = md mod n

= 63121 mod 247 = 232